# STEPC: A Pixel-wise Nonuniformity Correction Framework for Photon-Counting CT in Multi-material Imaging Scenarios

### v1

@@ Abstract
Photon-counting computed tomography (PCCT) has demonstrated significant advancements in recent years; however, pixel-wise detector response nonuniformity remains a key challenge, frequently manifesting as ring artifacts in reconstructed images. Existing correction methods exhibit limited generalizability in complex multi-material scenarios, such as contrast-enhanced imaging. This study introduces a Signal-to-Uniformity Error Polynomial Calibration (STEPC) framework to address this issue. STEPC first fits multi-energy projections using a 2D polynomial surface to generate ideal references, then applies a nonlinear multi-energy polynomial model to predict and correct pixel-wise nonuniformity errors. The model is calibrated using homogeneous slab phantoms of different materials, including PMMA, aluminum, and iodinated contrast agents, enabling correction for both non-contrast and contrast-enhanced imaging. Experiments were performed on a custom Micro-PCCT system with phantoms and mouse. Correction performance of STEPC was evaluated using the mean local standard deviation (MLSD) in the projection domain and the ring artifact deviation (RAD) on the reconstructed images. STEPC consistently outperformed existing correction methods in both non-contrast and contrast-enhanced scenarios. It achieved the lowest MLSD and RAD for both phantoms and mouse scans. These results indicate that STEPC provides a robust and practical solution for correcting detector nonuniformity in multi-material PCCT imaging, witch position it as a promising general-purpose calibration framework for photon-counting CT systems.
{IEEEkeywords}
photon counting CT, detector nonuniformity correction, polynomial calibration, ring artifact suppression, contrast-enhanced imaging.
1 Introduction
[…]
Photon-counting computed tomography (PCCT) has gained considerable attention [1, 2] due to its advantages in ultra-high-resolution imaging [2], beam hardening reduction [3], material decomposition [4], K-edge imaging [5] and more. However, PCCT still faces several challenges, including energy threshold bias and nonuniform pixel responses arising from variations in detector crystal and post-processing electronics. These effects degrade energy resolution and cause pixel-level response nonuniformity, often resulting in ring artifacts in the reconstructed images [6, 7, 8, 9].
@@ 1 Introduction
Various strategies have been proposed to suppress or correct pixel-wise nonuniformity in photon-counting CT, which can be broadly categorized into three main approaches: (1) Energy threshold calibration, which directly calibrates detector’s energy thresholds using monochromatic synchrotron sources, radioactive isotopes, X-ray fluorescence, or K-edge materials [10, 11]; (2) Ring artifact removal methods, which suppress artifacts in the projection or image domain through post-processing [12, 13, 14]; (3) Phantom-based measurement calibration, which uses calibration scans of known materials to learn a mapping from measured signals to ideal responses [15, 16, 17, 18]. Each approach has its own advantages and limitations. Energy threshold calibration can directly fix detector threshold bias but typically requires monochromatic X-ray sources and detector repositioning, witch is not readily available for most researchers. Ring artifact removal is simple to apply, but their correction capability is often limited, particularly in the presence of low-frequency concentric artifacts [7] or more severe nonlinear detector behaviors [19]. Phantom-based calibration directly models detector nonuniformity characteristics and is widely used as a key preprocessing step before reconstruction.
@@
Among phantom-based measurement calibration methods, traditional single-energy correction techniques [20, 15, 21, 22] can improve uniformity in single-material scenarios but perform poorly when multiple materials (e.g., soft tissue, bone, or iodine contrast agents) are present. Material decomposition methods [23, 24, 17] could address multi-material calibration, but their practicality is restricted by current photon-counting detectors (PCDs) limited energy thresholds (typically $\leq 2$), allowing only dual-material decomposition. This renders them inadequate for complex tasks, such as contrast-enhanced imaging. Additionally, most methods demand precise phantom positioning and thickness measurements, increasing operational complexity.
To overcome these challenges, we propose Signal-to-Uniformity Error Polynomial Calibration (STEPC), a novel framework for predicting and correcting photon-counting detector uniformity error responses. The method consists of four main steps. First, ideal reference projections are generated by applying a second-order 2D polynomial fit to flat-field projections acquired from combinations of calibration materials, avoiding the need for precise phantom alignment and measurement of material thickness or density. Then, The nonuniformity error is calculated by comparing the ideal projections with the actual measured projections. Next, A multi-energy polynomial regression model is fitted to learn the nonlinear relationship between the measured multi-energy projections and the corresponding nonuniformity errors. Finally, during actual imaging, the trained model is used to predict and correct detector nonuniformity in the acquired projections.
Unlike material decomposition approaches, the proposed method is robust to multi-material objects even under limited energy threshold numbers. By combined multi-material calibration slab phantoms (e.g., PMMA, Al, and contrast agents), the framework adapts to various imaging scenarios, including both non-contrast and contrast-enhanced imaging. The key contributions of this work include:
(1) A novel calibration framework for correcting PCDs nonuniformity in multi-material imaging scenarios, which is easy to implement and remains effective even with only two energy thresholds.
(2) Design of multi-material calibration slab phantoms, including PMMA, AL, and iodixanol, enabling effective correction for both general and contrast-enhanced imaging.
(3) Experimental validation on both phantoms and mouse confirms the necessity of material-specific calibration and demonstrates the superior performance of the proposed method in multi-material imaging scenarios.
2 Related Works
@@
Typically, different incident spectra are obtained using homogeneous phantoms with difference material (such as PMMA and AL) and thickness. The reference signal is commonly selected as either a single pixel value or the mean value [20, 16]. Alternative approaches include: (1) 1D polynomial fitting of projections to account for spatial variations in X-ray intensity or phantom thickness [25]. (2) directly using of measured thickness values as reference (common in material decomposition methods) [15, 18, 21, 22], and (3) simulation of ideal signals using known geometry, source spectra, and detector response functions [26, 27, 28].
@@
where $p$ is the polynomial order (typically 3), and the coefficients $c_{ij}$ and $d_{ij}$ are determined from calibration scans using difference thicknesses combinations of two basis materials.
@@
Among the above methods, FF and STC calibrate each energy bin independently, ignoring the additional information between energy bins. Moreover, FF does not consider object-induced changes in spectra. STC accounts for spectral changes but assumes a single material, which limits accuracy for biological tissues composed of multiple components (e.g., soft tissue and bone). Some methods extend single-energy correction to multi-material scenarios, such as Feng et al.[7] estimated the thickness of one material, then segmented other components using soft thresholding, but requiring additional steps such as segmentation, forward, and backward projections. Other approaches, PETC[21] transforms aluminum-equivalent thickness into PMMA-equivalent thickness using bin-wise projection models, but requires iterative per-pixel Gauss-Newton optimization (up to 1000 iterations), resulting in high computational cost. ATC allows for varying incident spectra but assumes a linear detector response, which limits its ability to handle the nonlinear characteristics of practical systems. PMDC improves correction by incorporating multi-material spectral changes and multi-energy bin information, but remains limited by the number of energy thresholds. Most commercial PCCT systems offer only two thresholds, witch allows only dual-material decomposition and makes it unsuitable for contrast-enhanced imaging. As shown in Figure 1, the incident spectra of non-contrast mouse scans can be modeled using PMMA and aluminum, whereas iodine-enhanced cases require iodine as an additional calibration material. Moreover, PMDC relies on accurate phantom thickness measurements and involves a signal-to-thickness conversion, which introduces additional uncertainty and may amplify noise.
3 Methods
Figure 2: Workflow of the proposed STEPC method. Step 1: Estimation of ideal projections for calibration slab phantoms; Step 2: Calculation of pixel-wise nonuniformity errors; Step 3: Fitting of the empirical multi-energy polynomial calibration model; Step 4: Projection correction for various imaging scenarios (red arrows indicate noticeable pixel response nonuniformity before correction).
@@ 3 Methods
To help explain the principle of the proposed method, the physical origin of the non-uniform detector response in photon-counting detectors (PCDs) is first described in this section. For the $j$th pixel in energy bin $E_{k}$, the measured detector counts can be expressed as:

### v2

# STEPC: A Multi-energy Nonuniform Response Calibration Framework for Photon-Counting Micro-CT in Multi-material Imaging

@@ Abstract
Photon-counting computed tomography has demonstrated significant advancements in recent years; however, micro photon-counting CT (Micro-PCCT) systems are still limited by pixel-wise detector response nonuniformity, which degrades measurement uniformity across detector pixels and commonly produces ring artifacts in reconstructed images. Existing calibration methods exhibit limited generalizability in complex multi-material scenarios, such as contrast-enhanced imaging. This study introduces a Signal-to-Nonuniformity Error Polynomial Calibration (STEPC) framework based on measurement nonuniformity error modeling to address this issue. STEPC first fits multi-energy projections using a 2D polynomial surface to generate ideal references, then applies a nonlinear multi-energy polynomial model to predict and correct pixel-wise nonuniformity errors. The model is calibrated using homogeneous slab phantoms of different materials, including PMMA, aluminum, and iodinated contrast agents, enabling correction for both non-contrast and contrast-enhanced imaging. Experiments were performed on a custom Micro-PCCT system with phantoms and mouse. Correction performance of STEPC was evaluated using the mean local standard deviation (MLSD) in the projection domain and the ring artifact deviation (RAD) on the reconstructed images. Compared with existing methods, STEPC achieved an average MLSD reduction of at least 21.58% and reduced RAD by at least 14.18%, consistently yielding the best performance in both non-contrast and contrast-enhanced scenarios. Furthermore, STEPC can be readily extended to compensate for beam hardening effects within the same calibration framework. Quantitative material decomposition results indicate that the proposed method preserves measurement accuracy across different basis materials. These results demonstrate that STEPC provides a robust and practical calibration framework that significantly improves detector measurement uniformity in Micro-PCCT systems for multi-material spectral imaging.
I Introduction
Photon-counting computed tomography (PCCT) has gained considerable attention due to its ability to directly measure photon counts in multiple energy windows, enabling ultra-high spatial resolution imaging, material decomposition, reduced beam hardening, and K-edge imaging[44, 35, 36, 8, 4]. Despite the numerous advantages of PCCT, its image quality critically depends on the uniformity of detector pixel responses. In practical systems, variations in detector crystals, electronic gain, and readout circuitry result in energy-dependent pixel response nonuniformity and threshold bias. These effects are particularly pronounced in Micro-PCCT systems, where reduced pixel sizes exacerbate pixel-to-pixel response variability. Such detector measurement nonuniformities propagate through the image reconstruction process and manifest as ring artifacts in the reconstructed images [52, 41, 10, 6, 49].
@@ I Introduction
Various strategies have been proposed to suppress or correct pixel-wise detector nonuniformity in photon-counting CT, which can be broadly categorized into three main approaches: (1) hardware calibration or control methods, (2) image processing-based methods, and (3) phantom-based measurement calibration. These approaches are often complementary and can be combined to achieve optimal correction performance. Hardware calibration focuses on the detector’s energy threshold calibration, which directly calibrates detector’s energy thresholds using monochromatic synchrotron sources, radioactive isotopes, X-ray fluorescence, or K-edge materials [45, 31]. However, this approach typically requires monochromatic sources and detector repositioning, which is not readily available for most researchers. Hardware control strategies mitigate nonuniformity by randomly moving the sample [16] or the detector [53], though this increases scan time and demands precise mechanical control. Helical scanning, widely used in clinical CT, can also convert ring-like artifacts to helix-like artifacts, thereby substantially reducing visible rings [28].
Image processing-based methods include projection-domain stripe suppression [27], pixel-wise gain correction [20, 3, 12], and ring removal in the reconstructed image by transforming it into the polar domain, performing stripe suppression, and then converting back to cartesian coordinates [48, 47, 50]. Dual-domain approaches have also been developed to enhance correction capability [40]. These methods are easy to implement and have been widely used; however, their correction capability is often limited, especially for low-frequency concentric artifacts [10] or severe nonlinear detector responses [43], and they may even introduce new artifacts in complex object structures [28]. Recently, deep learning has shown increasing promise for ring artifact suppression. Fang et al. [9] employed deep learning methods to remove ring artifacts by comparing models trained in the image domain, projection domain, and polar coordinate system, and further proposed a comprehensive model that effectively integrates multi-domain information for improved ring artifact suppression. Lv et al. [25] applied deep learning for simultaneous denoising and ring artifact correction in PCCT, while Yuan et al. [51] employed CNN and RNN architectures for ring artifact removal. Hein et al. [13] integrated a CNN with a spectral loss to improve correction performance in PCCT, and Sha et al. [38] proposed a dual-domain information fusion method using Transformer with a customized polar coordinate loss. Su et al. [39] designed a neural network framework that employs multilevel discrete wavelet transform (DWT) to achieve independent ring artifacts correction and interaction guidance of global and local features. Despite their effectiveness, these methods rely heavily on large paired training datasets, making practical deployment challenging. To address this, Hein et al. [14] further introduced a “Syn2Real” pipeline that efficiently synthesizes realistic ring artifacts directly in the image domain.
@@
Phantom-based measurement calibration, which uses calibration scans of known material phantoms to learn a mapping from measured signals to ideal responses [18, 29, 17, 22]. Compared with hardware-based methods, it is easier to implement, and unlike image processing approaches, it directly models detector nonuniformity characteristics from the system itself. Consequently, it serves as a key step in many CT reconstruction pipelines for effectively suppressing ring artifacts. Among phantom-based measurement calibration methods, traditional single-energy correction techniques [37, 18, 19, 11] can improve uniformity in single-material scenarios but perform poorly when multiple materials (e.g., soft tissue, bone, or iodine contrast agents) are present. Material decomposition methods [46, 2, 17] could address multi-material calibration, but their practicality is restricted by current photon-counting detectors (PCDs) limited energy thresholds (typically $\leq 2$), restricting them to dual-material decomposition in the projection domain. Although multi-material decomposition can be achieved by introducing additional constraints [23], such methods rely heavily on prior information and may introduce new biases. These limitations make existing techniques inadequate for complex tasks such as contrast-enhanced imaging. Furthermore, most of these methods require precise and time-consuming phantom calibration procedures. A recently proposed off-center water phantom calibration technique reduces this complexity [21]; however, it remains limited in handling multi-material scenarios.
In this study, we propose the Signal-to-Nonuniformity Error Polynomial Calibration (STEPC), a phantom-based calibration framework that characterizes and corrects material-dependent nonuniform response of PCDs in Micro-CT systems. STEPC models detector nonuniformity by attributing errors to intrinsic detector non-ideal factors and material-dependent X-ray attenuation. A multi-energy polynomial error model incorporates the attenuation characteristics of different materials to predict and correct pixel-wise measurement nonuniformity error. The effectiveness of this method is validated on a custom Micro-PCCT system using phantoms with different materials as well as mouse imaging with and without contrast agents. STEPC remains effective for multi-material imaging scenarios even with only two energy thresholds, and can also be readily extended to compensate for beam hardening. This approach provides a more accurate and comprehensive framework for modeling and calibrating photon-counting detector nonuniformity, and has the potential to serve as a general-purpose calibration framework for Micro-PCCT systems.
The rest of this article is organized as follows. Section II reviews related work on phantom-based calibration methods and introduces several baseline methods for comparison. Section  III presents the physical model of detector nonuniformity error and describes the steps of STEPC, along with its extension to beam hardening correction. Section IV details the experimental setup and evaluation metrics. Section V presents the main results from phantom and mouse imaging. Section VI discusses the advantages, limitations, and potential future directions of the proposed method. Finally, Section VII concludes the article.
II Related Works
@@
Typically, different incident spectra are obtained using homogeneous phantoms with different material (such as PMMA and AL) and thickness. The reference signal is commonly selected as either a single pixel value or the mean value [37, 29]. Alternative approaches include: (1) 1D polynomial fitting of projections to account for spatial variations in X-ray intensity or phantom thickness [1]. (2) directly using measured thickness values as reference (common in material decomposition methods) [18, 22, 19, 11], and (3) simulation of ideal signals using known geometry, source spectra, and detector response functions [7, 32, 24].
@@
where $p$ is the polynomial order (third-order in this study), and the coefficients $c_{ij}$ and $d_{ij}$ are determined from calibration scans using different thicknesses combinations of two basis materials.
@@
Among the above methods, FF and STC calibrate each energy bin independently, ignoring the additional information between energy bins. Moreover, FF does not consider object-induced changes in spectra. STC accounts for spectral changes but assumes a single material, which limits accuracy for biological tissues composed of multiple components (e.g., soft tissue and bone). Some methods extend single-energy correction to multi-material scenarios, such as Feng et al.[10] estimated the thickness of one material, then segmented other components using soft thresholding, but requiring additional steps such as segmentation, forward, and backward projections. Other approaches, PETC[19] transforms aluminum-equivalent thickness into PMMA-equivalent thickness using bin-wise projection models, but requires iterative per-pixel Gauss-Newton optimization (up to 1000 iterations), resulting in high computational cost. ATC allows for varying incident spectra but assumes a linear detector response, which limits its ability to handle the nonlinear characteristics of practical systems. PMDC improves correction by incorporating multi-material spectral changes and multi-energy bin information, but remains limited by the number of energy thresholds. Most commercial PCCT systems offer only two thresholds, which allows only dual-material decomposition and makes it unsuitable for contrast-enhanced imaging. As shown in Fig. 1, the incident spectra of non-contrast mouse scans can be modeled using PMMA and aluminum, whereas iodine-enhanced cases require iodine as an additional calibration material. Moreover, PMDC relies on accurate phantom thickness measurements and involves a signal-to-thickness conversion, which introduces additional uncertainty and may amplify noise.
III Methods
III-A Physical basis of nonuniform detector response
@@
To help explain the principle of the proposed method, the physical origin of the non-uniform detector response in photon-counting detectors (PCDs) is first described in this section. For the $j$th pixel in energy threshold $E_{k}$, the measured detector counts can be expressed as:
$N_{k,j}=\int_{E_{k}-\Delta E_{k,j}}^{E_{\max}}\int_{0}^{E_{\max}}S(E)e^{-\int_{l}{\mu(E,\vec{x})dl}}R_{j}(E^{\prime},E)\,dE\,dE^{\prime}$ (10)
where $\Delta E_{k,j}$ denotes the threshold shift of pixel $j$, $S(E)$ is the input X-ray spectrum, $E$ denotes the photon energy, $R_{j}(E^{\prime},E)$ represents the pixel-specific photon-counting detector response function, $E^{\prime}$ denotes the detected energy, $E_{\max}$ is the maximum photon energy determined by the tube voltage. Assuming the existence of an ideal or reference response function $R_{j}^{\mathrm{ideal}}(E^{\prime},E)$ that ensures uniformity across pixels, the corresponding ideal detector count is:
$N_{k,j}^{\mathrm{ideal}}=\int_{E_{k}}^{E_{\max}}\int_{0}^{E_{\max}}S(E)e^{-\int_{l}{\mu(E,\vec{x})dl}}R_{j}^{\mathrm{ideal}}(E^{\prime},E)\,dE\,dE^{\prime}$ (11)
Thus, the nonuniformity error is:
$\begin{split}&N_{k,j}^{\mathrm{error}}=N_{k,j}-N_{k,j}^{\mathrm{ideal}}\\ &\approx\int_{E_{k}}^{E_{\max}}\int_{0}^{E_{\max}}S(E)e^{-\int_{l}{\mu(E,\vec{x})dl}}\Delta R_{j}(E^{\prime},E)\,dE\,dE^{\prime}\\ &+\Delta E_{k,j}\int_{0}^{E_{\max}}S(E)\,e^{-\int_{l}\mu(E,\vec{x})\,dl}\,R_{j}^{\mathrm{ideal}}(E_{k},E)\,dE.\end{split}$ (12)
where $\Delta R_{j}(E^{\prime},E)=R_{j}(E^{\prime},E)-R_{j}^{\mathrm{ideal}}(E^{\prime},E)$. The detailed derivation of the approximation in Eq. (12) is provided in Appendix A. The above equation indicates that $N_{k,j}^{\mathrm{error}}$ arises from two primary sources: the pixel-dependent variation in detector response $\Delta R_{j}(E^{\prime},E)$ and the local energy threshold shift $\Delta E_{k,j}$. Both terms are modulated by the energy-dependent incident spectrum $S(E)e^{-\int_{l}\mu(E,\vec{x})\,dl}$, which depends on the material composition and attenuation properties along the X-ray path. Consequently, the nonuniformity error is a nonlinear functional jointly determined by the spectral characteristics of the object and the pixel-wise spectral response of the detector. Therefore, the nonuniformity estimated from single calibration material cannot be directly generalized to other materials or imaging conditions. We will also later prove that it is impossible to estimate the nonuniformity error for multi-material objects using only single-energy information.
III-B Signal-to-Nonuniformity Error Polynomial Calibration (STEPC)
Figure 2: Workflow of the proposed STEPC method. Step 1: Estimation of ideal projections for calibration slab phantoms; Step 2: Calculation of pixel-wise nonuniformity errors; Step 3: Generation of the calibration table by fitting the empirical multi-energy polynomial model; Step 4: Projection correction for various imaging scenarios (red arrows indicate noticeable pixel response nonuniformity before correction).
We propose a STEPC method that directly utilizes multi-energy projection information to predict and correct detector nonuniformity for multi-material imaging. The detailed workflow is illustrated in Fig. 2 and consists of the following steps:
Step 1: Ideal projection estimation. STEPC first measures the transmission through different combinations of multiple material slab phantoms. The log-transformed flat-field corrected projection data are then computed as $P(E_{k})=-\log(I_{E}^{\mathrm{FF}})$ to reduce the influence of X-ray source and detector instabilities. Inspired by Altunbas et al. [1], a second-order polynomial fit is applied to the 2D projection image for each energy bin $E_{k}$ to obtain the ideal projection:
$P_{s}^{\mathrm{ideal}}(x,y,E_{k})=\sum_{i+j\leq 2}b_{ij,s}^{(k)}x^{i}y^{j}$ (13)
Here, $(x,y)$ are the pixel coordinates, and $s$ indexes the calibration combinations of multiple material slabs. The coefficients $b_{ij,s}^{(k)}$ are obtained by minimizing the mean square error between $P_{s}(E_{k})$ and $P_{s}^{\mathrm{ideal}}(E_{k})$. We use MATLAB’s regress function [42] to solve for these coefficients.
This low-order polynomial approximation leverages the fact that both the X-ray beam intensity distribution and the transmitted signal through homogeneous or layered slab phantoms vary smoothly across the detector. The spatial dependence introduced by the source heel effect, beam filtration, and oblique incidence geometry changes gradually over the field of view, and thus can be effectively captured by a quadratic surface. As shown in Fig. 2(Step1), this polynomial fitting compensates for spatial variations in X-ray intensity and angular-dependent transmission thickness in the slab phantom, without requiring precise geometric measurements, thereby simplifying the calibration process.
$\Delta P_{s}^{\mathrm{ideal}}(x,y,E_{k})=P_{s}(x,y,E_{k})-P_{s}^{\mathrm{ideal}}(x,y,E_{k})$ (14)
As shown in Fig. 2(Step2), the nonuniformity error is dependent on the incident spectrum, which varies with material type and thickness. For single-material cases, single-energy projections can roughly estimate the error. However, for multi-materials, single-energy projections fail to capture the error accurately. In contrast, it illustrates that dual-energy information more effectively characterizes the error surface, allowing accurate modeling of errors caused by multi-materials.
Step 3: Calibration table fitting. As described in Eq. 12, the nonuniformity error arises from both detector-related non-ideal factors and material-dependent X-ray attenuation. By incorporating projections from all energy channels as input, the model captures spectral variations associated with the attenuation characteristics of different materials. An empirical polynomial model is then employed to implicitly entangle these effects within a unified formulation, enabling more accurate and robust correction in multi-material imaging scenarios. The model is formulated as follows:
$\hat{\Delta P_{s}}(x,y,E_{k})=\sum_{|\boldsymbol{\alpha}|\leq p}c_{\boldsymbol{\alpha}}^{(k)}(x,y)\prod_{n=1}^{N}P_{s}(x,y,E_{n})^{\alpha_{n}}$ (15)
where $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{N})$ is the multi-index vector, $|\boldsymbol{\alpha}|=\sum_{n=1}^{N}\alpha_{n}\leq p$ is the total polynomial degree (third-order in this study), $N$ is the number of energy thresholds, and $c_{\boldsymbol{\alpha}}^{(k)}(x,y)$ are the polynomial coefficients for energy channel $E_{k}$. These coefficients are determined by minimizing the mean square error of the estimator for ideal error $\Delta P_{s}^{\mathrm{ideal}}$:
$MSE=\mathbb{E}_{s}\left[\left(\hat{\Delta P_{s}}(x,y,E_{k})-\Delta P_{s}^{\mathrm{ideal}}(x,y,E_{k})\right)^{2}\right]$ (16)
Step 4: Correction. Finally, real object projection data are corrected by directly inputting the raw multi-energy projections into the fitted polynomial model to estimate the nonuniformity error:
$\hat{\Delta P}(x,y,E_{k})=\sum_{|\boldsymbol{\alpha}|\leq p}c_{\boldsymbol{\alpha}}^{(k)}(x,y)\prod_{n=1}^{N}P(x,y,E_{n})^{\alpha_{n}}$ (17)
and the corrected projection is obtained as:
$P^{\mathrm{corr}}(x,y,E_{k})=P(x,y,E_{k})-\hat{\Delta P}(x,y,E_{k})$ (18)
Notably, $c_{\boldsymbol{\alpha}}^{(k)}(x,y)$ indicates that the calibration and correction are independently performed for each detector pixel $(x,y)$ and each energy bin $E_{k}$.
Conventional methods such as STC and PMDC are primarily based on thickness estimation models, whose correction capability in multi-material imaging is inherently limited by the number of effective energy bins. ATC, on the other hand, compensates for count-domain biases induced by energy threshold variations; however, detector nonuniformity in photon-counting CT arises from multiple sources, and a simple affine transformation may be insufficient to capture the resulting nonlinear behavior. In contrast, STEPC directly formulates a nonuniformity error model in the projection domain, attributing projection inconsistencies to the combined effects of detector non-idealities (e.g., threshold variations and response differences) and multi-material X-ray attenuation. By incorporating projections from multiple energy bins, STEPC enriches multi-material attenuation information, while the polynomial model implicitly and effectively entangles detector non-idealities with material-dependent attenuation effects.
Furthermore, STEPC avoids the need for accurate thickness or density measurements of calibration phantoms by using 2D polynomial fitting to generate ideal reference projections. In addition, despite having only two energy thresholds, STEPC can incorporate iodine-based calibration to capture spectral variations from contrast agents, soft tissue, and bone, enabling effective correction in complex imaging scenarios.
III-C Extension to Beam Hardening Correction
The proposed STEPC framework can be further extended to correct for beam hardening (BH) effects. The overall procedure follows a similar structure to STEPC. In Step 1, the ideal projection $P_{s}^{\mathrm{ideal}}(x,y,E_{k})$ for each calibration slab combination $s$ is obtained as described previously. The difference lies in Step 2, where the beam hardening induced projection error is computed as:
$\begin{split}\Delta P_{s}^{\mathrm{BHideal}}&(x,y,E_{k})=\\ &P_{s}^{\mathrm{BHideal}}(x,y,E_{k})-P_{s}^{\mathrm{ideal}}(x,y,E_{k})\end{split}$ (19)
where the beam hardening affected ideal projection $P_{s}^{\mathrm{BHideal}}$ is modeled as:
$P_{s}^{\mathrm{BHideal}}(x,y,E_{k})=\sum_{m=1}^{M}T_{m}^{(s)}(x,y)\mu_{m}(E_{k})$ (20)
Here, $T_{m}^{(s)}$ denotes the thickness of the $m$th material in the $s$th slab combination, $M$ is the number of materials, and $\mu_{m}(E_{k})$ represents the effective linear attenuation coefficient at energy bin $E_{k}$. In biomedical imaging, beam hardening is mainly caused by low-density soft tissue and high-density bone, thus PMMA and aluminum ($M=2$) are typically employed for BH calibration. The $\mu_{m}(E_{k})$ values can be estimated via linear fitting from PMMA and aluminum slab measurements.
Using the multi-energy projection information, the beam-hardening correction term is then estimated with a polynomial regression model analogous to STEPC:
$\begin{split}\hat{\Delta P}_{s}^{\mathrm{BHcorr}}&(x,y,E_{k})=\\ &\sum_{|\boldsymbol{\alpha}|\leq p}c_{\mathrm{BHcorr},\boldsymbol{\alpha}}^{(k)}(x,y)\prod_{n=1}^{N}P_{s}^{\mathrm{ideal}}(x,y,E_{n})^{\alpha_{n}}\end{split}$ (21)
Here, the polynomial degree $p$ is also set to three, and the coefficients $c_{\mathrm{BHcorr},\boldsymbol{\alpha}}^{(k)}(x,y)$ are determined by minimizing the mean square error over all calibration combinations $s$:
$\begin{split}&MSE=\\ &\mathbb{E}_{s}\left[\left(\hat{\Delta P}_{s}^{\mathrm{BHcorr}}(x,y,E_{k})-\Delta P_{s}^{\mathrm{BHideal}}(x,y,E_{k})\right)^{2}\right]\end{split}$ (22)
Finally, beam hardening correction is applied to the nonuniformity corrected projection $P^{\mathrm{corr}}(x,y,E_{k})$:
$\begin{split}\hat{\Delta P}^{\mathrm{BHcorr}}&(x,y,E_{k})=\\ &\sum_{|\boldsymbol{\alpha}|\leq p}c_{\mathrm{BHcorr},\boldsymbol{\alpha}}^{(k)}(x,y)\prod_{n=1}^{N}P^{\mathrm{corr}}(x,y,E_{n})^{\alpha_{n}}\end{split}$ (23)
and the beam hardening corrected projection is obtained as:
$P^{\mathrm{BHcorr}}(x,y,E_{k})=P^{\mathrm{corr}}(x,y,E_{k})-\hat{\Delta P}^{\mathrm{BHcorr}}(x,y,E_{k})$ (24)
It should be noted that the beam hardening correction extension is not inherently coupled with the detector nonuniformity model. It can be applied independently as a standalone correction module and still provides effective beam hardening mitigation. Its applicability to other calibration methods is further demonstrated in Section V-F.
Figure 3: Schematic of the system geometry and multi-material slab phantoms. (a) Geometry configuration of the calibration setup; (b) three types of slab phantoms: PMMA, aluminum, and iodixanol; (c) schematic of iodixanol phantom container; (d) combinations of PMMA and aluminum slabs with varying thicknesses; (e) combinations of iodixanol solution slabs with PMMA and aluminum slabs.
IV Experiments
IV-A Micro Photon Counting CT and Calibration Materials